uestion6 0.25 pts with interest at 10%, what is the benefit-cost ratio for this government project? nitial Cost Additional costs at the end of year 1 and year 2 Benefits at end of year 1 and year so/year 2 Annual benefits at end of year 3 $101015/year through year 10 $203180 $29800/year Enter your answer as follow: 12.34
Answer #1

Present worth of benefit=101015((1/1.1)^3+(1/1.1)^4+...+(1/1.1)^10)=101015*4.76=480783.3

Present worth of Cost=203180+29800((1/1.1)^1+(1/1.1)^2))=203180+29800*(1.74)=254899

Benefit-Cost Ratio=PW of Benefit/ PW of Cost=480783.3/254899=1.886=1.89

Benefit Cost Ratio =1.89